OpenStudy (anonymous):

Chem: Calculate the moleseeded to prepare each a. 1.0 L of a 3.0M NaCl solution b. 0.40 L of a 1.0M KBr solution c. 125 mL of a 2.0 M MgCl2 solution

OpenStudy (anonymous):

the correct answers are a. 3 moles b. .40 moles c. .25 moles but i don't know how to get them!

OpenStudy (anonymous):

a. 3.0M=3moles/liter(1liiter)=3moles b. 1.0M=1moles/liter(.4liter)=.4 moles b. 2.0M=2moles/liter(.125liters)=.25moles

OpenStudy (anonymous):

the 3.0M numbers give you how many moles there are per L, so if you multiply by how many liters there are you get the amount of moles

OpenStudy (anonymous):

Just to clarify on what physopholy said. You need to cancel L to get the moles: \[1.0 L * 3.0 M = 1.0 L * 3.0 \frac{moles}{L}=3.0 moles NaCL\] \[0.40L*1.0M=0.40L*1.0\frac{moles}{L}=0.40 moles KBr\] \[125ml*2.0M=0.125L*2.0\frac{moles}{L}=0.25 moles MgCl_2\] Note you have to keep the correct Significant figures which is 2 in each case.